define angle of deviation in a prism

Grade:9

3 Answers

Prajwal kr
11 years ago

It the angle through which light ray deviates from the original straight line path.

sourav singh
11 years ago

the angle made by path of the incident ray with the path of the refracted ray is the angle of deviation.

Akash Kumar Dutta
11 years ago

Dear Koushik,
the difference between the angle of incident ray and emer gent ray is angle of deviation.
the formula for it is A+D=i+e ...where A=angle of prism,D=deviation,i=incident ray angle,e=emergent ray angle.
Also in an inclined plane the deviation changes.
Regards.